OpenStudy (anonymous):

Solve the system by graphing. y=-2x+2 5x+y=5

OpenStudy (saifoo.khan):

substitute the value of the first equation into second one.

OpenStudy (saifoo.khan):

5x+(-2x+2) = 5 Solve, you will get a value for x.

hero (hero):

y = -2x+2 y = 5-5x y = y -2x+2 = 5-5x 3x = 3 x =1 y =0

OpenStudy (anonymous):

now how do i go about graphing it? is it only one point

hero (hero):

Yes, I gave you the way to graph them by solving both for y

OpenStudy (anonymous):

oh ok thanks

hero (hero):

Notice the y equals

hero (hero):

You will see that they intersect at the point (1,0)

OpenStudy (saifoo.khan):

for the first one, -2 is the slope and 2 is the y intercept.

OpenStudy (anonymous):

yes y=y. so the point is (1,0) and just graph that

hero (hero):

Using a graphing calculator would be a good idea

OpenStudy (anonymous):

I dont have one that is the problem lol

hero (hero):

I'm sorry to hear that

OpenStudy (anonymous):

its ok